# On the Ability of Deep Networks to Learn Symmetries from Data: A Neural Kernel Theory

### v1

@@ Acknowledgments
Appendix A Derivation of the spectral error
@@
$(y_{1},\dots,y_{N},y_{t})\sim\mathcal{N}\left(0,\begin{bmatrix}K&k(\vec{x},x_{t})\\ k(\vec{x},x_{t})^{T}&k(x_{t},x_{t})\end{bmatrix}\right),$ (20)
@@
$\mu_{t|\mathcal{D}}$ $=(K^{-1}k(\vec{x},x_{t}))^{T}\vec{y},$ (21)
$\sigma_{t|\mathcal{D}}$ $=k(x_{t},x_{t})-(K^{-1}k(\vec{x},x_{t}))^{T}k(\vec{x},x_{t}),$ (22)
@@
In Section 3.2, we take the Gaussian (RBF) kernel as the kernel for Gaussian Process regression. In later sections, we consider various deep network architectures in the infinite width limit. These also behave like Gaussian Processes, with a kernel that is specified by their architecture (Neal, 1996b, ; Jacot et al.,, 2018). Our proof for the spectral formula of the generalization error does not depend on the specific kernel used, as long as the kernel matrix is circulant over a group of points generated by the action of a cyclic group. We demonstrate elsewhere that this is the case both for the Gaussian kernel, as well as all the deep neural kernels that we analyze in this paper (MLP and ConvNets).
We are now ready to present the derivation of Eq. 7. We start by presenting a depiction of the conditioning procedure for the case in which the training set consists of a single point (Fig. 8(a)). In the following, we denote labels as $y_{0}$ and $y_{1}$, and their respective values as $\mu_{0}$ and $\mu_{1}$. In practice, the $y$s are the training set labels, and the $\mu$s are their values. In this case, the training label is $y_{0}$, while the test label is $y_{1}$. The two labels are jointly distributed according to a 2 dimensional Gaussian, with mean 0 and covariance matrix $K$. This matrix depends on the values of $x_{0},x_{1}$, and the kernel function $k$. We represent the distribution as an ellipse, which is understood as a level set of the covariance matrix $K$, centered at the mean, 0, in a way similar to the usual representation of confidence levels of a multivariate Gaussian distribution. Conditioning over the known training label $y_{0}$ can be interpreted geometrically as slicing the ellipse with a vertical line at coordinate $\mu_{0}$. We thus obtain a segment (i.e., a 1-dimensional ellipse), the midpoint of which is the point $(\mu_{0},\mu_{1|0})$, where $\mu_{1|0}$ denotes the conditional value of $y_{1}$ given $x_{0},x_{1}$ and the value of $y_{0}$. As a consequence, the result of GP regression can thus be interpreted geometrically as finding the $y_{1}$ coordinate of the center of the slice. Doing this will reward us with Eq. 7.
@@
We can make this geometric task easier by turning the ellipse into a circle; we can do this by performing a whitening operation on this vector space. Such whitening is effectively obtained by the inverse of the square root of the covariance matrix $K$. Denoting the transformed values by priming them, our task is now to find $\vec{\mu}_{1}^{\prime}$ given $\vec{\mu}_{0}^{\prime}$ [^6] We have now shifted to a vector notation for these values as, unlike in the original space, both components can be nonzero.. It is clear that the segment produced by the conditioning turns now into a chord of the circle. This makes the whitened vector $\vec{\mu}^{\prime}$, which is defined so that it points to the midpoint of the segment, perpendicular to the segment itself. This means we can write the following geometric condition:
@@
$M_{ij}$ $=\mu_{i}\delta_{ij},$ (28)
$E_{ij}$ $=\frac{1}{\sqrt{2N}}\omega^{ij},$ (29)
$E^{-1}_{ij}$ $=\frac{1}{\sqrt{2N}}\omega^{-ij},$ (30)
$\Lambda^{-1/2}_{ij}$ $=\frac{1}{\sqrt{\lambda_{i}}}\delta_{ij}.$ (31)
@@
$M^{\prime}_{ji}$ $=(ZM)_{ji}$ (32)
$=\sum_{k}Z_{jk}M_{ki}$ (33)
$=\sum_{k}(E\Lambda^{-1}E^{-1})_{jk}M_{ki}$ (34)
$=\sum_{klm}E_{jl}\Lambda^{-1/2}_{lm}E^{-1}_{mk}M_{ki}$ (35)
$=\sum_{klm}\frac{1}{\sqrt{2N}}\omega^{jl}\frac{1}{\sqrt{\lambda_{l}}}\delta_{lm}\frac{1}{\sqrt{2N}}\omega^{-mk}\mu_{k}\delta_{ki}$ (36)
$=\sum_{l}\frac{1}{2N}\omega^{jl}\frac{1}{\sqrt{\lambda_{l}}}\omega^{-li}\mu_{i}$ (37)
$=\sum_{l}\frac{1}{2N}\frac{\mu_{i}\omega^{l(j-i)}}{\sqrt{\lambda_{l}}}.$ (38)
@@
$\sum_{j}M_{ji}^{\prime}M_{j0}^{\prime}$ $=\sum_{j}(\sum_{l}\frac{1}{2N}\frac{\mu_{i}\omega^{l(j-i)}}{\sqrt{\lambda_{l}}})(\sum_{k}\frac{1}{2N}\frac{\mu_{0}\omega^{kj}}{\sqrt{\lambda_{k}}})$ (40)
$=\sum_{lkj}\frac{1}{(2N)^{2}}\mu_{i}\mu_{0}\frac{\omega^{l(j-i)}\omega^{kj}}{\sqrt{\lambda_{l}\lambda_{k}}}$ (41)
$=\sum_{lkj}\frac{1}{2N}\frac{\mu_{i}\mu_{0}\omega^{-il}}{\sqrt{\lambda_{l}\lambda_{k}}}\frac{1}{2N}\omega^{j(l+k)}.$ (42)
@@
$\sum_{j}\frac{1}{2N}\omega^{j(l+k)}=\delta_{l,2N-k}.$ (43)
@@
$\sum_{j}M_{ji}^{\prime}M_{j0}^{\prime}$ $=\sum_{kl}\frac{1}{2N}\frac{\mu_{i}\mu_{0}\omega^{-il}}{\sqrt{\lambda_{l}\lambda_{k}}}\delta_{l,2N-k}$ (44)
$=\sum_{k}\frac{1}{2N}\frac{\mu_{i}\mu_{0}\omega^{-i(2N-k)}}{\sqrt{\lambda_{2N-k}\lambda_{k}}}$ (45)
$=\sum_{k}\frac{1}{2N}\frac{\mu_{i}\mu_{0}\omega^{ik}}{\lambda_{k}}.$ (46)
@@
$\sum_{i=0}^{2N-1}\sum_{k}\frac{1}{2N}\frac{\mu_{i}\omega^{ik}}{\lambda_{k}}=0.$ (47)
@@
Note how, for $i=0$ (that is, the missing point) we recover the unknown value $\mu_{0}$ for which we want to solve. With this substitution, we get
$\sum_{i=0}^{2N-1}\frac{1}{2N}(\mu\omega^{iN}+(\mu_{0}-\mu)\delta_{i0})\sum_{k}\frac{\omega^{ik}}{\lambda_{k}}=0.$ (49)
@@
$\sum_{i=0}^{2N-1}\frac{\mu}{N}\sum_{k}\frac{\omega^{i(k+N)}}{\lambda_{k}}+\sum_{i=0}^{2N-1}\frac{\mu_{0}}{2N}\sum_{k}\frac{\omega^{ik}}{\lambda_{k}}\delta_{i0}-\sum_{i=0}^{2N-1}\frac{\mu}{2N}\delta_{i0}\sum_{k}\frac{\omega^{ik}}{\lambda_{k}}=0$ (50)
@@
    $\sum_{i=0}^{2N-1}\sum_{k}\frac{\mu}{2N}\frac{\omega^{i(k+N)}}{\lambda_{k}}$ $=\mu\sum_{k}\frac{1}{\lambda_{k}}\sum_{i=0}^{2N-1}\frac{1}{2N}\omega^{i(k+N)}$ (51)
    $=\mu\sum_{k}\frac{1}{\lambda_{k}}\delta_{k,-N}$ (52)
    $=\mu\frac{1}{\lambda_{N}}.$ (53)
@@
    $\mu_{0}\sum_{k}\frac{1}{2N}\frac{1}{\lambda_{k}}=\mu_{0}\langle\lambda^{-1}\rangle,$ (54)
@@
    $\mu\langle\lambda^{-1}\rangle.$ (55)
@@
$\mu\frac{1}{\lambda_{N}}+\mu_{0}\langle\lambda^{-1}\rangle-\mu\langle\lambda^{-1}\rangle=0,$ (56)
@@
The formula for the spectral error, then, assuming $\mu=1$, is
$\varepsilon_{s}=\frac{\lambda_{N}^{-1}}{\langle\lambda^{-1}\rangle}.$ (58)
A.1 Extension to multiple missing points
In the case of multiple missing points, we can find an equivalent of Eq. 58. It requires solving an $p$ dimensional linear system, where $p$ is the number of missing points.
@@ A.1 Extension to multiple missing points
$\sum_{m\in[p]}\mu_{m}\frac{1}{2N}\sum_{l}\frac{\omega^{l(j-m)}}{\lambda_{l}}=\mu\left(\frac{1}{2N}\sum_{n\in[p]}\omega^{nN}\sum_{k}\frac{\omega^{k(j-n)}}{\lambda_{k}}-\frac{\omega^{jN}}{\lambda_{N}}\right).$ (59)
@@
$A_{jm}=\frac{1}{2N}\sum_{l}\frac{\omega^{l(j-m)}}{\lambda_{l}},\quad b_{j}=\mu\left(\frac{1}{2N}\sum_{n\in[p]}\omega^{nN}\sum_{k}\frac{\omega^{k(j-n)}}{\lambda_{k}}-\frac{\omega^{jN}}{\lambda_{N}}\right).$ (61)

### v2

@@ 5 Discussion
Appendix A A (Very Brief) Introduction to Neural Kernel Theories
Here we briefly introduce the mathematical connection that exists between deep neural networks and kernel methods. We reuse and modify the words of Jan Gerken (https://gapindnns.github.io/_pages/wide_nns.html), with permission:
The number of neurons in each layer of a neural network is called the width $N$ of that layer. When considering the limit in which the width of all hidden layers goes to infinity (assuming that the weights of the network scale as $1/\sqrt{N}$ to avoid activity divergence), the neural network simplifies dramatically. By an argument using the central limit theorem, one can show that in the infinite width limit, the neurons follow a collective Gaussian distribution known as a Gaussian process. Intuitively, the fluctuations from all the neurons cancel out. This effect was known for a long time as the Neural Network Gaussian Process (NNGP) theory (Neal, 1996) and characterizes the neural network at initialization, i.e. before training has begun. Furthermore, the neural kernel—characterizing dot product similarity of data points in the last infinitely-wide layer of the network—becomes independent of the specific initialization chosen and can be computed using recursive layer-by-layer expressions.
In 2018, a seminal paper by Jacot et al. (2018) showed that the simplifications go even further than that: In the infinite width limit (again assuming that the weights scale as $1/\sqrt{N}$), the dynamics remain manageable even during training. Specifically, they proved that the kernel of an infinitely wide network remains constant throughout training. This kernel, called the Neural Tangent Kernel (NTK), is equal to the NNGP kernel with additional corrective terms. Briefly, this simplification is possible because all weights in the inner layers remain in the vicinity of their initial values during training, allowing to linearize the network around the initial condition. Under the simple but realistic training paradigm of gradient descent of the mean-squared-error loss, the training dynamics can then be solved analytically in closed form and the prediction of the trained network on any input be computed. The output of the trained network is again a Gaussian process.
These simplifications in the infinite width limit give powerful insights into the behavior of neural networks.
Appendix B Derivation of the Spectral Error
@@
$(y_{1},\dots,y_{N},y_{t})\sim\mathcal{N}\left(0,\begin{bmatrix}K&k(\vec{x},x_{t})\\ k(\vec{x},x_{t})^{T}&k(x_{t},x_{t})\end{bmatrix}\right),$
@@
$\mu_{t|\mathcal{D}}$ $=(K^{-1}k(\vec{x},x_{t}))^{T}\vec{y},$
$\sigma_{t|\mathcal{D}}$ $=k(x_{t},x_{t})-(K^{-1}k(\vec{x},x_{t}))^{T}k(\vec{x},x_{t}),$
@@
In Section 4.3, we take the Gaussian (RBF) kernel as the kernel for Gaussian Process regression. In later sections, we consider various deep network architectures in the infinite width limit. These also behave like Gaussian Processes, with a kernel that is specified by their architecture (Neal, 1996; Jacot et al., 2018).
Our proof for the spectral formula of the generalization error does not depend on the specific kernel used, as long as the kernel (Gram) matrix is circulant over a group of points generated by the action of a cyclic group. We demonstrate elsewhere that this is the case both for the Gaussian kernel, as well as all the deep neural kernels that we analyze in this paper (MLP and ConvNets).
We are now ready to present the derivation of Eq. 1. This equation describes the most likely error in prediction resulting from Gaussian Process regression on a dataset generated by cyclic group action. We refer to this most likely error as the result of kernel regression in main text. We do not study the fluctuations of errors around this mean.
We start by presenting a depiction of the conditioning procedure for the case in which the training set consists of a single point (Fig. 8(a)). In the following, we denote labels as $y_{0}$ and $y_{1}$, and their respective values as $\mu_{0}$ and $\mu_{1}$. In practice, the $y$s are the training set labels, and the $\mu$s are their values. In this case, the training label is $y_{0}$, while the test label is $y_{1}$. The two labels are jointly distributed according to a 2 dimensional Gaussian, with mean 0 and covariance matrix $K$. This matrix depends on the values of $x_{0},x_{1}$, and the kernel function $k$. We represent the distribution as an ellipse, which is understood as a level set of the covariance matrix $K$, centered at the mean, 0, in a way similar to the usual representation of confidence levels of a multivariate Gaussian distribution. Conditioning over the known training label $y_{0}$ can be interpreted geometrically as slicing the ellipse with a vertical line at coordinate $\mu_{0}$. We thus obtain a segment (i.e., a 1-dimensional ellipse), the midpoint of which is the point $(\mu_{0},\mu_{1|0})$, where $\mu_{1|0}$ denotes the conditional value of $y_{1}$ given $x_{0},x_{1}$ and the value of $y_{0}$. As a consequence, the result of GP regression can thus be interpreted geometrically as finding the $y_{1}$ coordinate of the center of the slice. Doing this will reward us with Eq. 1.
@@
We can make this geometric task easier by turning the ellipse into a circle; we can do this by performing a whitening operation on this vector space. Such whitening is effectively obtained by the inverse of the square root of the covariance matrix $K$. Denoting the transformed values by priming them, our task is now to find $\vec{\mu}_{1}^{\prime}$ given $\vec{\mu}_{0}^{\prime}$. [^10] We have now shifted to a vector notation for these values as, unlike in the original space, both components can be nonzero. It is clear that the segment produced by the conditioning turns now into a chord of the circle. This makes the whitened vector $\vec{\mu}^{\prime}$, which is defined so that it points to the midpoint of the segment, perpendicular to the segment itself. This means we can write the following geometric condition:
@@
$M_{ij}$ $=\mu_{i}\delta_{ij},$
$E_{ij}$ $=\frac{1}{\sqrt{2N}}\omega^{ij},$
$E^{-1}_{ij}$ $=\frac{1}{\sqrt{2N}}\omega^{-ij},$
$\Lambda^{-1/2}_{ij}$ $=\frac{1}{\sqrt{\lambda_{i}}}\delta_{ij}.$
@@
$M^{\prime}_{ji}$ $=(ZM)_{ji}$
$=\sum_{k}Z_{jk}M_{ki}$
$=\sum_{k}(E\Lambda^{-1}E^{-1})_{jk}M_{ki}$
$=\sum_{klm}E_{jl}\Lambda^{-1/2}_{lm}E^{-1}_{mk}M_{ki}$
$=\sum_{klm}\frac{1}{\sqrt{2N}}\omega^{jl}\frac{1}{\sqrt{\lambda_{l}}}\delta_{lm}\frac{1}{\sqrt{2N}}\omega^{-mk}\mu_{k}\delta_{ki}$
$=\sum_{l}\frac{1}{2N}\omega^{jl}\frac{1}{\sqrt{\lambda_{l}}}\omega^{-li}\mu_{i}$
$=\sum_{l}\frac{1}{2N}\frac{\mu_{i}\omega^{l(j-i)}}{\sqrt{\lambda_{l}}}.$
@@
$\sum_{j}M_{ji}^{\prime}M_{j0}^{\prime}$ $=\sum_{j}(\sum_{l}\frac{1}{2N}\frac{\mu_{i}\omega^{l(j-i)}}{\sqrt{\lambda_{l}}})(\sum_{k}\frac{1}{2N}\frac{\mu_{0}\omega^{kj}}{\sqrt{\lambda_{k}}})$
$=\sum_{lkj}\frac{1}{(2N)^{2}}\mu_{i}\mu_{0}\frac{\omega^{l(j-i)}\omega^{kj}}{\sqrt{\lambda_{l}\lambda_{k}}}$
$=\sum_{lkj}\frac{1}{2N}\frac{\mu_{i}\mu_{0}\omega^{-il}}{\sqrt{\lambda_{l}\lambda_{k}}}\frac{1}{2N}\omega^{j(l+k)}.$
@@
$\sum_{j}\frac{1}{2N}\omega^{j(l+k)}=\delta_{l,2N-k}.$
@@
$\sum_{j}M_{ji}^{\prime}M_{j0}^{\prime}$ $=\sum_{kl}\frac{1}{2N}\frac{\mu_{i}\mu_{0}\omega^{-il}}{\sqrt{\lambda_{l}\lambda_{k}}}\delta_{l,2N-k}$
$=\sum_{k}\frac{1}{2N}\frac{\mu_{i}\mu_{0}\omega^{-i(2N-k)}}{\sqrt{\lambda_{2N-k}\lambda_{k}}}$
$=\sum_{k}\frac{1}{2N}\frac{\mu_{i}\mu_{0}\omega^{ik}}{\lambda_{k}}.$
@@
$\sum_{i=0}^{2N-1}\sum_{k}\frac{1}{2N}\frac{\mu_{i}\omega^{ik}}{\lambda_{k}}=0.$
@@
Note how, for $i=0$ (that is, the missing point) we recover the unknown value $\mu_{0}$ for which we want to solve.
With this substitution, we get
$\sum_{i=0}^{2N-1}\frac{1}{2N}(\mu\omega^{iN}+(\mu_{0}-\mu)\delta_{i0})\sum_{k}\frac{\omega^{ik}}{\lambda_{k}}=0.$
@@
$\sum_{i=0}^{2N-1}\frac{\mu}{N}\sum_{k}\frac{\omega^{i(k+N)}}{\lambda_{k}}+\sum_{i=0}^{2N-1}\frac{\mu_{0}}{2N}\sum_{k}\frac{\omega^{ik}}{\lambda_{k}}\delta_{i0}-\sum_{i=0}^{2N-1}\frac{\mu}{2N}\delta_{i0}\sum_{k}\frac{\omega^{ik}}{\lambda_{k}}=0$
@@
    $\sum_{i=0}^{2N-1}\sum_{k}\frac{\mu}{2N}\frac{\omega^{i(k+N)}}{\lambda_{k}}$ $=\mu\sum_{k}\frac{1}{\lambda_{k}}\sum_{i=0}^{2N-1}\frac{1}{2N}\omega^{i(k+N)}$
    $=\mu\sum_{k}\frac{1}{\lambda_{k}}\delta_{k,-N}$
    $=\mu\frac{1}{\lambda_{N}}.$
@@
    $\mu_{0}\sum_{k}\frac{1}{2N}\frac{1}{\lambda_{k}}=\mu_{0}\langle\lambda^{-1}\rangle,$
@@
    $\mu\langle\lambda^{-1}\rangle.$
@@
$\mu\frac{1}{\lambda_{N}}+\mu_{0}\langle\lambda^{-1}\rangle-\mu\langle\lambda^{-1}\rangle=0,$
@@
The formula for the spectral error, then, is
$\varepsilon_{s}\triangleq\frac{\mu_{0}-\mu}{\mu}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{=\frac{\lambda_{N}^{-1}}{\langle\lambda^{-1}\rangle}}.$ (8)
B.1 Extension to Multiple Missing Points
In the case of multiple missing points, we can find an equivalent of Eq. 8. It requires solving a $p$ dimensional linear system, where $p$ is the number of missing points.
@@ B.1 Extension to Multiple Missing Points
$\sum_{m\in[p]}\mu_{m}\frac{1}{2N}\sum_{l}\frac{\omega^{l(j-m)}}{\lambda_{l}}=\mu\left(\frac{1}{2N}\sum_{n\in[p]}\omega^{nN}\sum_{k}\frac{\omega^{k(j-n)}}{\lambda_{k}}-\frac{\omega^{jN}}{\lambda_{N}}\right).$
@@
$A_{jm}=\frac{1}{2N}\sum_{l}\frac{\omega^{l(j-m)}}{\lambda_{l}},\quad b_{j}=\mu\left(\frac{1}{2N}\sum_{n\in[p]}\omega^{nN}\sum_{k}\frac{\omega^{k(j-n)}}{\lambda_{k}}-\frac{\omega^{jN}}{\lambda_{N}}\right).$